Sarpedon

Homer and Chapman

Compiled, November 28, 2009

From Chapman’s Homer’s Iliad, Book XII, lines 290330:

Nor had great Hector and his friends the rampire overrun, If heav’n’s great counsellor, high Jove, had not inflam’d his son Sarpedon (like the forest’s king when he on oxen flies) Against the Grecians: his round targe he to his arm applies, Brass-leav’d without, and all within, thick ox-hides quilted hard, The verge nail’d round with rods of gold; and with two darts prepar’d He leads his people: as ye see a mountain-lion fare, Long kept from prey; in forcing which, his high mind makes him

dare Assault upon the whole full fold, though guarded never so With well-arm’d men and eager dogs; away he will not go, But venture on, and either snatch a prey, or be a prey. So far’d divine Sarpedon’s mind, resolv’d to force his way Through all the fore-fights, and the wall: yet since he did not see Others as great as he in name, as great in mind as he, He spake to Glaucus: ‘Glaucus, say, why are we honour’d more Than other men of Lycia in place, with greater store Of meats and cups, with goodlier roofs, delightsome gardens, walks, More lands and better, so much wealth that court and country talks Of us and our possessions, and every way we go, Gaze on us as we were their gods? This where we dwell is so: The shores of Xanthus ring of this, and shall we not exceed As much in noise? Come, be we great in deed As well as look; shine not in gold, but in the flames of fight, That so our neat-arm’d Lycians may say: “See, these are right Our kings, our rulers; these deserve to eat and drink the best; These govern not ingloriously: these thus exceed the rest, Do more than they command to do.” O friend, if keeping back Would keep back age from us, and death, and that we might not

wrack In this life’s human sea at all, but that deferring now We shunn’d death ever, nor would I half this vain valour show, Nor glorify a folly so, to wish thee to advance: But since we must go, though not here, and that, besides the chance Propos’d now, there are infinite fates of other sort in death, Which (neither to be fled nor ’scap’d) a man must sink beneath, Come, try we if this sort be ours: and either render thus Glory to others, or make them resign the like to us.’ This motion Glaucon shifted not, but (without words) obey’d; Foreright went both, a mighty troop of Lycians followed.